# Tracing the impacts of Mount Pinatubo eruption on global climate using spatially-varying changepoint detection

### v1

@@ 3.1 Changepoint detection and estimation
Let $\boldsymbol{\gamma}_{0R}$ and $\boldsymbol{\gamma}_{1R}$ be the vector $\gamma_{0R}({\bf s}_{i})$ and ${\gamma_{1R}({\bf s}_{i})}$ for $i=1,\ldots,N$, respectively. Let $\mathbf{U}$ be the vector stacking up all $U({\bf s}_{i},t_{j})$ for $i=1,\ldots,N$ and $j=1,\ldots,M$, and let $\log(\boldsymbol{\Delta})$ be the vector of element-wise logarithm of $\boldsymbol{\Delta}$. In our next level of the hierarchical model, we provide priors for the latent processes in model (1).
@@
For computational simplicity, we assume that $\boldsymbol{\Sigma_{U}}$ has a separable space-time covariance structure with exponential covariance functions in both space and time, i.e. $\boldsymbol{\Sigma_{U}}=\sigma^{2}_{U}\mathbf{R}(\phi_{U})\otimes\mathbf{R}(\psi_{U}),$ where the $(i,j)$th element of $\mathbf{R}(\theta)$ is $R(\theta)_{ij}=\exp(-\theta h)$ for a distance $h$ either in space or time, and $\otimes$ is the Kronecker product. Since our data spans the entire globe, we employed the great circle distance as the distance metric for the spatial correlation matrix. While a positive definite covariance function defined on Euclidean space may not necessarily be valid on spheres, the exponential covariance function has been shown to be valid on spheres when the Euclidean distance is replaced with the great circle distance. (Gneiting, 2013; Huang et al., 2011). We assume $\mathbf{\Sigma}_{\gamma_{0}}=\sigma^{2}_{\gamma_{0}}\mathbf{R}(\psi_{\gamma_{0}})$, $\mathbf{\Sigma}_{\gamma_{1}}=\sigma^{2}_{\gamma_{1}}\mathbf{R}(\psi_{\gamma_{1}})$ and $\boldsymbol{\Sigma}_{\Delta}=\sigma^{2}_{\Delta}\mathbf{R}(\psi_{\Delta})$, where $\mathbf{R}(\psi_{\gamma_{0}})$, $\mathbf{R}(\psi_{\gamma_{1}})$ and $\mathbf{R}(\psi_{\Delta})$ are also governed by an exponential covariance function with great circle distance.
@@ 3.2 MCMC Sampling
where $\circ$ denotes the Schur product operator, and $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\boldsymbol{\epsilon}_{1},\boldsymbol{\epsilon}_{2}$ denote the $MN$-length vector of $\mu_{1}({\bf s}_{i},t_{j})$, $\mu_{2}({\bf s}_{i},t_{j})$, $\epsilon_{1}({\bf s}_{i},t_{j})$ and $\epsilon_{2}({\bf s}_{i},t_{j})$ for $i=1,\ldots,N$ and $j=1,\ldots,M$. Let $\boldsymbol{\tau}$ be the vector of $\begin{pmatrix}\tau({\bf s}_{1}),\ldots,\tau({\bf s}_{N})^{T}\end{pmatrix}$. The full conditional likelihood of $\bf Y$, $f(\mathbf{Y}\lvert\mathbf{U},\boldsymbol{\tau},\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\sigma_{1}^{2},\sigma_{2}^{2})$, is
@@
There are two main challenges in MCMC sampling for our model. First, the Appendix shows that obtaining the posterior for $\mathbf{U}$ involves inverting the covariance matrix ${\bf\Sigma}_{Y}^{-1}+{\bf\Sigma}_{U}^{-1}$. Although $\bf\Sigma_{U}$ is separable, (8) shows that the matrix ${\bf\Sigma}_{Y}$ involves vectors $\mathbf{1}_{\tau}^{-}$ and $\mathbf{1}_{\tau}^{+}$, which depends on $\tau({\bf s})$ and thus cannot be separated into a Kronecker product of $M\times M$ and $N\times N$ matrices. This significantly amplifies the computational cost associated with sampling $\mathbf{U}$. Secondly, sampling both $\mathbf{U}$ and $\boldsymbol{\mu}_{\tau}$ can in practice lead to an identifiability issue, since both $\mathbf{U}$ and $\boldsymbol{\mu}_{\tau}$ have the potential to capture the trend in the data. While integrating out $\mathbf{U}$ from the model could solve these problems, retaining $\mathbf{U}$ is more desirable from a computational perspective. This stems from the fact that $\mathbf{Y}$ is no longer conditionally independent when the conditioning on $\mathbf{U}$ is removed, as indicated by the non-diagonal covariance matrix ${\bf\Sigma}_{Y}+{\bf\Sigma}_{U}$ in (9). Consequently, sampling other parameters whose full conditionals involve the likelihood of $\mathbf{Y}$ becomes significantly more intricate.
To address the first issue, we propose a simple but effective approach. We assume, without the loss of generality, that the variance increases after the changepoint and write the post-changepoint measurement error as $\boldsymbol{\epsilon}_{2}=\boldsymbol{\epsilon}_{1}+\boldsymbol{\epsilon}_{\gamma},$ where $\boldsymbol{\epsilon}_{\gamma}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}N(0,\sigma^{2}_{\gamma})$ is a vector of white noise independent of $\boldsymbol{\epsilon}_{1}.$ Then, we can rewrite $f(\boldsymbol{Y}\lvert\cdot)$ in (8) as
@@
To address the identifiability between $\bf U$ and $\boldsymbol{\mu}_{\tau}$ in sampling, we adopt the Partially Collapsed Gibbs (PCG) sampler proposed by Van Dyk and Park (2008), which allows us to remove the conditioning on $\mathbf{U}$ when sampling $\boldsymbol{\tau}$ and $\boldsymbol{\alpha}=(\alpha_{0},\gamma_{0},\gamma_{1})^{T}$. We refer to Van Dyk and Park (2008) and Van Dyk and Jiao (2015) for more details on how to derive the PCG sampler from an ordinary Gibbs sampler. Figure 3 illustrates the sampling procedure for our model under the PCG sampler. It is important to note that, unlike ordinary Gibbs samplers, the order of draws in the PCG sampler must be maintained, as permuting the order may alter the stationary distribution of the chain in the PCG sampler (Van Dyk and Jiao, 2015).
@@
The following discusses the sampling for each parameter in turn. The full conditionals for $\mathbf{U}$ and $\boldsymbol{\epsilon_{\gamma}}$ are $MN$-dimensional multivariate normal distributions. To sample $\mathbf{U}$, we leverage the compatibility of a Kronecker product plus a constant diagonal matrix with the eigenvalue decomposition (Stegle et al., 2011), which allows us to sample $\mathbf{U}$ using eigendecompositions of $M\times M$ and $N\times N$ matrices instead of having to factorize a $MN\times MN$ covariance matrix. While the posterior for $\boldsymbol{\epsilon_{\gamma}}$ is no longer identically distributed, the Appendix shows that its independence structure is preserved. Thus, posterior samples can be drawn easily even when the spatial and/or temporal dimension is high. See Appendix for more details.
@@
The parameters $\sigma^{2}_{U},\sigma_{1}^{2},\sigma_{2}^{2},\sigma^{2}_{\Delta},\sigma^{2}_{\gamma_{0}},\sigma^{2}_{\gamma_{1}},\boldsymbol{\beta},$ and $\boldsymbol{\alpha}=(\alpha_{0},\gamma_{0},\gamma_{1})^{T}$ can be assigned conjugate priors and estimated via a Gibbs sampler. We specify weakly independent normal prior for $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ and inverse gamma priors for variance parameters $\sigma^{2}_{1},\sigma^{2}_{2},\sigma^{2}_{U},$ and $\sigma^{2}_{\Delta}.$ For $\sigma^{2}_{\gamma_{0}}$ and $\sigma^{2}_{\gamma_{1}},$ we use a truncated inverse gamma prior and choose the upper truncation points to be a small fraction of $\gamma_{0F}$ and $\gamma_{1F}.$ This encourages the magnitude of random effect to be smaller than that of the fixed effect, which further ensures identifiability. The parameters $\psi_{U},\phi_{U},\psi_{\Delta},\psi_{\gamma_{0}},\psi_{\gamma_{1}}$ are given uniform priors and are sampled via the Metropolis Hastings algorithm. For $\tau_{0},$ any discrete probability distribution that takes values in $\{1,2,\ldots,M\}$ can be used as a prior. The full conditional distributions and sampling details are outlined in the Appendix.
@@ 4 Simulation Study
We perform simulation studies to evaluate the accuracy of our model in detecting three different types of changes – mean shift only, variance shift only, and mean and variance shift. We also explore how the magnitude of change affects the model’s performance.
@@ 4.1 Setup
To generate data for simulation, we select $N=121$ spatial locations on an $11\times 1$ 11 grid and randomly choose one of the grid locations as $\mathbf{s_{0}}$. At each location, we consider a time series of length $M=61$ and generate $\tau(\mathbf{s})$ using (5), with $\tau_{0}=18,\boldsymbol{\beta}=(1.5,1)^{T},$ $\sigma^{2}_{\Delta}=1,$ and $\psi_{\Delta}=0.5$. These values were chosen such that the number of locations with changepoint |$|\{{\bf s}:\tau(\mathbf{s})<M\}|$ is comparable to the number of locations without changepoints |$|\{{\bf s}:\tau(\mathbf{s})\geq M\}|$ on average. Since a changepoint detection is equivalent to a classification problem, having a balance between the two classes ensures our evaluation focuses on method performance without being affected by class imbalance. Without loss of generality, we fix $\alpha_{0}=0$ and $\sigma^{2}_{1}=1$. We let $\gamma_{0}$ take values in $\{0,1.5,2,3\}$ and $\sigma^{2}_{2}$ in $\{3,5,10\}$ to study the model performance under different strengths of mean and variance shift. The parameters $\sigma^{2}_{U},$ $\phi_{U},$ and $\psi_{U}$ are set to 1, 1.5, and 2, respectively, to emulate spatial and temporal dependence observed in real data based on our exploratory data analysis of global surface temperature data. We use model (1) to generate data and run 100 simulations for each setting. The ratio of number of locations without changepoints to those with changepoints in our simulated data ranges from 32:89 to 101:20, with an average of 57:64.
@@
In implementing our method, we assume no prior knowledge about $\tau_{0}$ and use a discrete uniform prior. This is slightly more challenging than our real problem, since for the Pinatubo data, we know $\tau_{0}$ is likely right after the eruption. We give weakly informative prior for other unknown parameters. We use the posterior mode of $\min(M,\tau({\bf s}))$ as the estimated changepoint for location ${\bf s}$. If the mode for $\min(M,\tau({\bf s}))$ is at $M$, we determine that there is no changepoint at ${\bf s}$.
@@
To evaluate model performance, we use the metrics of false positive rate (FPR) and false negative rate (FNR) to measure the accuracy of the changepoint detection, and use root mean squared error (RMSE), empirical coverage of the 95% credible interval (CI), and length of the CI to assess the changepoint estimation. FPR is defined as the ratio of the number of falsely detected locations to the total number of locations that have changepoints detected, and FNR is the ratio of the number of falsely undetected locations to the total number of locations without changepoints detected. When calculating RMSE, we treat the true changepoint as having value M (i.e., $\tau(\mathbf{s})=M$) when there is no changepoint at location ${\bf s}$. A combination of a narrower credible interval with empirical coverage closer to the nominal level indicates a more precise uncertainty quantification.
We compare the FPR, FNR and RMSE of our spatial-temporal model with the univariate time series changepoint detection method offered by the R package changepoint. The latter method is applied to each of the spatial locations separately (hereinafter 1D method). We use cpt.mean, cpt.var, and cpt.meanvar functions for each type of change, all using modified Bayesian information criteria penalty for estimating at most one changepoint (Hinkley, 1970).

### v2

@@ 3.1 Changepoint detection and estimation
Let $\boldsymbol{\gamma}_{0R}$ and $\boldsymbol{\gamma}_{1R}$ be the vector of $\gamma_{0R}({\bf s}_{i})$ and $\gamma_{1R}({\bf s}_{i})$} for $i=1,\ldots,N$, respectively. Let $\mathbf{U}$ be the vector stacking up all $U({\bf s}_{i},t_{j})$ for $i=1,\ldots,N$ and $j=1,\ldots,M$, and let $\log(\boldsymbol{\Delta})$ be the vector of element-wise logarithms of $\boldsymbol{\Delta}$. In our next level of the hierarchical model, we provide priors for the latent processes in model (1).
@@
For computational simplicity, we assume that $\boldsymbol{\Sigma_{U}}$ has a separable space-time covariance structure with an exponential covariance function in both space and time, i.e. $\boldsymbol{\Sigma_{U}}=\sigma^{2}_{U}\mathbf{R}(\phi_{U})\otimes\mathbf{R}(\psi_{U}),$ where the $(i,j)$th element of $\mathbf{R}(\theta)$ is $R(\theta)_{ij}=\exp(-\theta h)$ for a distance $h$ either in space or time, and $\otimes$ is the Kronecker product. Since our data spans the entire globe, we employed the great circle distance as the distance metric for the spatial correlation matrix. While a positive definite covariance function defined on Euclidean space may not necessarily be valid on spheres, the exponential covariance function has been shown to be valid on spheres when the Euclidean distance is replaced with the great circle distance. (Gneiting, 2013; Huang et al., 2011). We assume $\mathbf{\Sigma}_{\gamma_{0}}=\sigma^{2}_{\gamma_{0}}\mathbf{R}(\psi_{\gamma_{0}})$, $\mathbf{\Sigma}_{\gamma_{1}}=\sigma^{2}_{\gamma_{1}}\mathbf{R}(\psi_{\gamma_{1}})$ and $\boldsymbol{\Sigma}_{\Delta}=\sigma^{2}_{\Delta}\mathbf{R}(\psi_{\Delta})$, where $\mathbf{R}(\psi_{\gamma_{0}})$, $\mathbf{R}(\psi_{\gamma_{1}})$ and $\mathbf{R}(\psi_{\Delta})$ are also governed by an exponential covariance function with great circle distance.
@@ 3.2 MCMC Sampling
where $\circ$ denotes the Schur product operator, and $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\boldsymbol{\epsilon}_{1},\boldsymbol{\epsilon}_{2}$ denote the $MN$-length vector of $\mu_{1}({\bf s}_{i},t_{j})$, $\mu_{2}({\bf s}_{i},t_{j})$, $\epsilon_{1}({\bf s}_{i},t_{j})$ and $\epsilon_{2}({\bf s}_{i},t_{j})$ for $i=1,\ldots,N$ and $j=1,\ldots,M$. Let $\boldsymbol{\tau}$ be the vector of $\begin{pmatrix}\tau({\bf s}_{1}),\ldots,\tau({\bf s}_{N})^{T}\end{pmatrix}$. The full conditional likelihood of $\bf Y$ given $\mathbf{U},\boldsymbol{\tau},\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ is
@@
There are two main challenges in MCMC sampling for our model. First, Section 1 of the Supplement (Shi-Jun et al., 2023) shows that obtaining the posterior for $\mathbf{U}$ involves inverting the covariance matrix ${\bf\Sigma}_{Y}^{-1}+{\bf\Sigma}_{U}^{-1}$. If ${\bf\Sigma}_{Y}$ is a constant diagonal matrix, then because $\bf\Sigma_{U}$ is separable we can sample $\mathbf{U}$ using eigendecompositions of $M\times M$ and $N\times N$ matrices instead of having to factorize a $MN\times MN$ covariance matrix (Stegle et al., 2011). However, (8) shows that the matrix ${\bf\Sigma}_{Y}$ involves vectors $\mathbf{1}_{\tau}^{-}$ and $\mathbf{1}_{\tau}^{+}$ which depends on $\tau({\bf s})$, and thus is not a constant diagonal matrix. As a consequence, we have to invert a $MN\times MN$ covariance matrix. This significantly amplifies the computational cost associated with sampling $\mathbf{U}$. Secondly, sampling both $\mathbf{U}$ and $\boldsymbol{\mu}_{\tau}$ can in practice lead to an identifiability issue, since both $\mathbf{U}$ and $\boldsymbol{\mu}_{\tau}$ have the potential to capture the trend in the data. While integrating out $\mathbf{U}$ from the model could solve these problems, retaining $\mathbf{U}$ is more desirable from a computational perspective. This stems from the fact that $\mathbf{Y}$ is no longer conditionally independent when the conditioning on $\mathbf{U}$ is removed, as indicated by the non-diagonal covariance matrix ${\bf\Sigma}_{Y}+{\bf\Sigma}_{U}$ in (9). Consequently, sampling other parameters whose full conditionals involve the likelihood of $\mathbf{Y}$ becomes significantly more intricate.
[…]
@@
To address the potential challenge of identifiability between $\bf U$ and $\boldsymbol{\mu}_{\tau}$ in sampling, we adopt the Partially Collapsed Gibbs (PCG) sampler proposed by Van Dyk and Park (2008), which allows us to remove the conditioning on $\mathbf{U}$ when sampling $\boldsymbol{\tau}$ and $\boldsymbol{\alpha}=(\alpha_{0},\gamma_{0},\gamma_{1})^{T}$. We refer to Van Dyk and Park (2008) and Van Dyk and Jiao (2015) for more details on how to derive the PCG sampler from an ordinary Gibbs sampler. Figure 3 illustrates the sampling procedure for our model under the PCG sampler. It is important to note that, unlike ordinary Gibbs samplers, the order of draws in the PCG sampler must be maintained, as permuting the order may alter the stationary distribution of the chain in the PCG sampler (Van Dyk and Jiao, 2015).
@@
The following discusses the sampling for each parameter in turn. The full conditionals for $\mathbf{U}$ and $\boldsymbol{\epsilon_{\gamma}}$ are $MN$-dimensional multivariate normal distributions. To sample $\mathbf{U}$, we leverage the compatibility of a Kronecker product plus a constant diagonal matrix with the eigenvalue decomposition (Stegle et al., 2011), which allows us to sample $\mathbf{U}$ using eigendecompositions of $M\times M$ and $N\times N$ matrices. While the posterior for $\boldsymbol{\epsilon_{\gamma}}$ is no longer identically distributed, Section 1 of the Supplement (Shi-Jun et al., 2023) shows that its independence structure is still preserved. Thus, posterior samples can be drawn easily even when the spatial and/or temporal dimension is high. See the Supplement Material (Shi-Jun et al., 2023) for more details.
@@
Finally, the parameters $\sigma^{2}_{U},\sigma_{1}^{2},\sigma_{2}^{2},\sigma^{2}_{\Delta},\sigma^{2}_{\gamma_{0}},\sigma^{2}_{\gamma_{1}},\boldsymbol{\beta},$ and $\boldsymbol{\alpha}=(\alpha_{0},\gamma_{0},\gamma_{1})^{T}$ are assigned conjugate priors and estimated via a Gibbs sampler. We specify weakly independent normal prior for $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ and inverse gamma priors for variance parameters $\sigma^{2}_{1},\sigma^{2}_{2},\sigma^{2}_{U},$ and $\sigma^{2}_{\Delta}.$ For $\sigma^{2}_{\gamma_{0}}$ and $\sigma^{2}_{\gamma_{1}},$ we use a truncated inverse gamma prior and choose the upper truncation points to be a small fraction of $\gamma_{0F}$ and $\gamma_{1F}.$ This encourages the magnitude of random effect to be smaller than that of the fixed effect, which further ensures identifiability. The parameters $\psi_{U},\phi_{U},\psi_{\Delta},\psi_{\gamma_{0}},\psi_{\gamma_{1}}$ are given uniform priors and are sampled via the Metropolis Hastings algorithm. For $\tau_{0},$ any discrete probability distribution that takes values in $\{1,2,\ldots,M\}$ can be used as a prior. The full conditional distributions and sampling details are outlined in Section 1 of the Supplement (Shi-Jun et al., 2023).
@@ 4 Simulation Study
We perform simulation studies to evaluate the accuracy of our model in detecting and estimating three different types of changes – mean shift only, variance shift only, and mean and variance shift. We compare our method with a univariate time series changepoint detection method derived by only considering $N=1$ in our model. This univariate method still considers temporal correlation but no spatial correlation and is applied to each of the spatial locations separately. We also explore how the magnitude of change affects the model’s performance.
@@ 4.1 Setup
To generate data for simulation, we select $N=121$ spatial locations on an $11\times 11$ grid and randomly choose one of the grid locations as $\mathbf{s_{0}}$. The grid is on a surface of a sphere with latitudes from $0^{\circ}$ to $50^{\circ}$ and longitudes from $0^{\circ}$ to $50^{\circ},$ both in an increment of $5^{\circ}$. At each location, we consider a time series of length $M=61$ and generate $\tau(\mathbf{s})$ using (5), with $\tau_{0}=18,\boldsymbol{\beta}=(1.5,1)^{T},$ $\sigma^{2}_{\Delta}=1,$ and $\psi_{\Delta}=0.5$. These values were chosen such that the number of locations with changepoint $|\{{\bf s}:\tau(\mathbf{s})<M\}|$ is comparable to the number of locations without changepoints $|\{{\bf s}:\tau(\mathbf{s})\geq M\}|$ on average. Since a changepoint detection is equivalent to a classification problem, having a balance between the two classes ensures our evaluation focuses on method performance without being affected by class imbalance.
We use model (1) to generate data but replace $\mu_{2}({\bf s},t)$ in model (3) by $\mu_{2}({\bf s},t)=\alpha_{0}+\gamma_{0}$. This simple version of $\mu_{2}({\bf s},t)$ serves the purpose of changepoint detection evaluation and offers a fair comparison with the classic method for which the default model contains no slope after the changepoint. Without loss of generality, we fix the pre-changepoint mean $\alpha_{0}=0$ and variance $\sigma^{2}_{1}=1$. We let $\gamma_{0}$ take values in $\{0,1.5,2,3\}$ and $\sigma^{2}_{\gamma}$ in $\{0,3,5\}$ to study the model performance under different strengths of mean and variance shift. The parameters $\sigma^{2}_{U},$ $\phi_{U},$ and $\psi_{U}$ are set to 1, 1.5, and 2, respectively, to emulate spatial and temporal dependence observed in real data based on our exploratory data analysis of global surface temperature data. We run 100 simulations for each setting. The ratio of the number of locations without changepoints to those with changepoints in our simulated data ranges from 32:89 to 101:20, with an average of 57:64.
@@
In implementing our method, we assume no prior knowledge about $\tau_{0}$ and use a discrete uniform prior. This is slightly more challenging than our real problem, since for the Pinatubo data, we know $\tau_{0}$ is likely right after the eruption. We give weakly informative prior for other unknown parameters. We use the posterior mode of $\min\{M,\tau({\bf s})\}$ as the estimated changepoint for location $\bf s$}. If the mode for $\min\{M,\tau({\bf s})\}$ is at $M$, we determine that there is no changepoint at $\bf s$}.
@@
To evaluate the performance of our method, we use the metrics of false positive rate (FPR) and false negative rate (FNR) to measure the accuracy of the changepoint detection, and use root mean squared error (RMSE), empirical coverage of the 95% credible interval (CI), and length of the CI to assess the changepoint estimation. FPR is defined as the ratio of the number of falsely detected locations to the total number of locations without changepoints, and FNR is the ratio of the number of falsely undetected locations to the total number of locations with changepoints. When calculating RMSE, we treat the true changepoint as having value $M$ (i.e., $\tau(\mathbf{s})=M$) when there is no changepoint at location $\bf s$}. A combination of a narrower credible interval with empirical coverage closer to the nominal level indicates a more precise uncertainty quantification. The formula for FPR, FNR, and RMSE are given below, where $\tau({\bf s})$ denotes the true changepoint for location $\bf s$}, $\hat{\tau}({\bf s})$ denotes the detected changepoint, and $\lvert A\rvert$ is the cardinality of set A:
$FPR$ $=\frac{FP}{FP+TN}=\frac{\lvert\{{\bf s}:\hat{\tau}({\bf s})<M,\tau({\bf s})=M\}\rvert}{\lvert\{{\bf s}:\tau({\bf s})=M\}\rvert}$
$FNR$ $=\frac{FN}{FN+TP}=\frac{\lvert\{{\bf s}:\hat{\tau}({\bf s})=M,\tau({\bf s})<M\}\rvert}{\lvert\{{\bf s}:\tau({\bf s})<M\}\rvert}$
$RMSE$ $=\sqrt{\frac{\sum_{\bf s}(\tau({\bf s})-\hat{\tau}({\bf s}))^{2}}{N}}$
We compare the FPR, FNR and RMSE of our spatio-temporal model with the univariate method applied to each location (hereinafter 1D method). Through experimentation, we picked a conservatively large iteration number of $20,000$ and a burn-in size of $10,000$ to ensure MCMC convergence based on Gelman-Rubin diagnostic (Gelman and Rubin, 1992).